OpenStudy (anonymous):

How do you solve 3x-10=x+30 ?

OpenStudy (hoblos):

subtract x from both side 2x -10 = 30 add 10 for both sides 2x = 40 divide both sides by 2 x = 20

OpenStudy (anonymous):

thankyou, this really helped, do you use this method to solve all equations?

OpenStudy (ash2326):

We have \[3x-10=x+30\] Subtract x from both sides \[3x-10-x=x+30-x\] \[3x-10- x=\cancel x+30-\cancel x\] \[2x-10=30\] Add 10 to both sides \[2x-10+10=30+10\] \[2x-\cancel{10}+\cancel {10}=30+10\] We get \[2x=40\] divide both sides by 2 \[\frac{2x}{2}=\frac{40}{2}\] \[x=20\] \we get \